Let X be a discrete random variable that follows a Poisson distribution with λ=4.

What is P(X<5|X>3)?

please answer to at least 3 decimal places
Answer #1

TOPIC:Poisson distribution.=> Given that, [ X ~ Poisson (x = 4 So, the pref of xis- 1 P(x+x) = { e. 4 ; x = 0, 1, 2, - ; otherwise. we need to find x P

7 Again, we have, I Law of = 1-P (x 23). I complementation] = 1- [P(x=0) + P (X=1) + P(x= 2) + p (x = 3) (ory the ball =--Let